OpenStudy (anonymous):

"find each variable. leave your answers in simplest radical form." can anybody help?

OpenStudy (anonymous):



OpenStudy (anonymous):

is it 7?

OpenStudy (anonymous):

if it is 7, x^2+7^2=10^2 x^2+49=100 x^2=51 x=square root of 51

OpenStudy (anonymous):

i got \[2\sqrt{14}\]

OpenStudy (anonymous):